# Towards Understanding Depth Perception in Foveated Rendering

### v1

@@ Abstract.
The true vision for real-time virtual and augmented reality is reproducing our visual reality in its entirety on immersive displays. To this end, foveated rendering leverages the limitations of spatial acuity in human peripheral vision to allocate computational resources to the fovea while reducing quality in the periphery. Such methods are often derived from studies on the spatial resolution of the human visual system and its ability to perceive blur in the periphery, enabling the potential for high spatial quality in real-time. However, the effects of blur on other visual cues that depend on luminance contrast, such as depth, remain largely unexplored. It is critical to understand this interplay, as accurate depth representation is a fundamental aspect of visual realism. In this paper, we present the first evaluation exploring the effects of foveated rendering on stereoscopic depth perception. We design a psychovisual experiment to quantitatively study the effects of peripheral blur on depth perception. Our analysis demonstrates that stereoscopic acuity remains unaffected (or even improves) by high levels of peripheral blur. Based on our studies, we derive a simple perceptual model that determines the amount of foveation that does not affect stereoacuity. Furthermore, we analyze the model in the context of common foveation practices reported in literature. The findings indicate that foveated rendering does not impact stereoscopic depth perception, and stereoacuity remains unaffected up to 2×2$\times$2 × stronger foveation than commonly used. Finally, we conduct a validation experiment and show that our findings hold for complex natural stimuli.
@@ 2.2. Depth Perception
To navigate 3D space, humans rely on a number of different depth cues, each being suitable and dominant in specific scenarios. For example, while we mostly rely on stereopsis (binocular vision), convergence, and accommodation for close objects, we primarily use the cue of aerial perspective for very distant targets (Cutting and Vishton, 1995). With the advent of stereoscopic displays, the recreation of binocular vision became feasible. This enabled the display of stereoscopic depth cues, such as stereopsis. Stereopsis is created by perceiving a difference in vergence angles between a focus point $p_{f}$ and a reference point $p_{r}$, which naturally occurs when those points are at different depths with respect to the eyes
@@
(Figure 2). The vergence angles can be described by $D(p)=\angle LpR$, denoting the angles between the two locations of the eyes in space ($L$ and $R$ ), and the respective points $p_{f}$ and $p_{r}$. Stereopsis is quantified by the disparity $d$ [], denoting the absolute difference of these vergence angles:
(1) $d=|D(p_{1})-D(p_{2})|.$
In case both vergence angles are the same, $d=0$ and no disparity is present. This can happen when both points are at the same position in space, the image is rendered with an interpupillary distance (IPD) of 0mm, or both points lie on the horopter (Blakemore, 1970).
@@ 3. Methods
The stimulus design minimizes the influence of all depth cues except binocular disparity. These cues include for example motion parallax, shading, texture density, and deformation. Monocular observation of the stimulus does not convey depth information, except for minor texture deformation. However, the visibility of the texture deformation is well below the resolvable disparity limit and, therefore, negligible (Bradshaw and Rogers, 1999).
@@ 3.1.1. Texture
50% Random Dot Patterns are widely used in vision science, particularly for researching depth perception (Julesz, 1960). They serve as a simple baseline pattern for our metric, exhibiting controllable features such as scale, frequency range, and dynamic range. As we aim to measure the lowest possible disparity threshold, we scale the random dot pattern to the optimally perceptible size. This scale equals the peak sensitivity value of the contrast sensitivity function (CSF) for the given eccentricity $\theta$. The stelaCSF [^1] https://github.com/gfxdisp/stelaCSF, retrieved 18. Oct 2024 (Mantiuk et al., 2022) yields a peak sensitivity $S_{F}(\theta)$ of $S_{F}(0^{\circ})=4.1$cpd, $S_{F}(10^{\circ})=1.8$cpd, and $S_{F}(0^{\circ})=1.3$cpd. We scale the dot pattern to match $S_{F}(\theta)$, by scaling the dots such that two dots equal one cycle.
@@
Additionally, the ring must be of a sufficient minimum width in the visual field to ensure good visibility. However, to reduce crosstalk between eccentricities and ensure accurate measurements, the width is minimized. To account for the varying sensitivity per eccentricity, we scale the ring wider with higher eccentricities. We set the ring’s width based on the CSF, choosing a width of 13 dots for both peripheral stimuli. This is approximately equal to $3.8^{\circ}$ in the visual field for $10^{\circ}$, and $5^{\circ}$ for $20^{\circ}$ eccentricity. For $0^{\circ}$ eccentricity we set a width of $6.7^{\circ}$, so that always 2 full cycles are visible. Additionally, to attenuate the depth cue of geometric deformation, the ring has a randomly dithered border (Figure 3).
@@ 3.1.2. Depth
eye’s image, while leaving it untouched in the left eye’s image ($p_{L}$). This shift introduces an intersection point $p_{I}$ which our brain perceives as being in front of the zero-plane (Figure 4). The disparity is measured between the point $p_{Z}$, which serves as a zero point reference (trough, $I_{D}(p_{Z})=0$), and $p_{I}$, the artificially created illusion point (peak, $I_{D}(p_{I})=1$). We generate the right eye’s image $I_{R}$ by resampling the left eye’s image $I_{L}$ at the respective image coordinates:
@@ 3.2.3. Participants
The first part of the study ($10^{\circ}$ and $20^{\circ}$) was conducted on 11 voluntary participants (1F, 10M), with an average age of 28 ± 4.7 years. All participants had normal or corrected-to-normal vision and were tested negative for stereo blindness. Nine participants reported prior experience in conducting perceptual experiments. The second part was conducted on 10 of the 11 previous participants (1F, 9M; mean age 28.1 ± 4.9 years). Eight participants reported prior experience. With the exception of the two participating authors, all participants were naive to the project.
@@ 4. Results and Perceptual Model
Figure 7. Plot of estimated thresholds per eccentricity, as well as the polynomial fits. The area of each blue point denotes the weight $w$ of the corresponding threshold. The orange scatter plots denote the mean and standard error per $\sigma$. The line plots show the retrieved parabolas, fitted through the weighted threshold data.
@@ 4.1. Threshold Estimation
Since the PEST procedure can lead to sparse sampling of important regions and random answers can render data unreliable, the Weibull fit can get unstable. We obtain the final threshold $T$ as well as its standard deviation $T_{\sigma}$ by bootstrapping the data 100 times. This also yields a relative uncertainty of $u=\frac{T_{\sigma}}{T}$ per $T$. Measurements with $u>0.3$ are treated as outliers and excluded from further evaluation, as we consider the PEST procedure to be too unstable and therefore unreliable. This affects ~20% of our measurements.
@@ 4.2. Perceptual Model
To estimate the parameters’ interpolation characteristics for varying $\theta$, we examine the characteristics of our fitted parabolas (Figure 7): The minimum threshold, associated with $p_{2}$, shifts to a higher $\sigma$ with increasing eccentricity. This indicates that the HVS tolerates higher blur levels in the periphery. Furthermore, the overall threshold $T$ shifts to higher values with increasing eccentricity, which resembles the y-offset encoded in $p_{3}$. The slope $p_{1}$ remains mostly unaltered. Figure 8 displays the parameter-values $p_{i}$ per eccentricity.
[…]
As expected, they resemble the intuitive understanding discussed above. To model the parabolas for a continuous $\theta$, we fit an exponential curve of the form $p_{i}(\theta)=ae^{b\theta}+c$. An exponential curves increase strictly, which resembles our data. The fitting yields:
@@
This model holds for our measured scope of $\sigma\in[0,15]$, and $\theta\in[0,20]$. However, we note that $p_{1}$ exhibits a steep slope for higher eccentricities, resulting in extreme values for $\theta>20$. If such conditions are expected we suggest replacing $p_{1}$ with a linear fit of $p_{1}=0.0034$.
@@ 5.1. Methods
We show participants half-split images of the natural scenes. The central part is blacked out up to $15^{\circ}$ eccentricity. We validate our model on high eccentricities, as the fovea remains almost unaffected by our model, yielding no measurable changes. As we allow the gaze to move within a margin of 3–5 from the fixation point, the effective blacking is ~10. Furthermore, an examination of $>15^{\circ}$ eccentricity retains enough binocular image space, as the headset only supports binocular vision up to $25^{\circ}$ eccentricity. The left and right halves of the image display different disparity levels: one half is rendered with 0mm IPD (no disparity), while the other half is rendered with a varying IPD (0–20mm) to control the overall magnitude of disparities in the scene (see Supplementary Material for vergence angle histograms). The IPD, and thus the overall scene depth, was adjusted in a staircase procedure. Participants were shown both scenes in full-resolution (Org) and foveated (Fov) rendering styles. Each participant first viewed the Forest scene (F), and then the Kitchen scene (K). The order of presented rendering styles was randomized, and the style sequence was counterbalanced across scenes (e.g. F-Fov, F-Org, K-Org, K-Fov). Before threshold estimation began, participants were allowed to freely explore each scene using their central vision to familiarize themselves with prominent depth cues. This step was added to reduce measurement noise caused by initial variability in scene familiarity, as identified in pilot testing. Participants were asked to identify the side of the image which exhibited disparity. To estimate the IPD threshold, the displayed disparity was adjusted using the ”Best PEST” procedure, as described in Section 3.2.1. The experiment was conducted using the same apparatus as described in Section 3.2.2. We conducted the study on 15 participants (13M, 1F, 1X; mean age 26.1 ± 2.5 years). 11 participants reported prior experience with perceptual studies. All had normal or corrected-to-normal vision.
@@ 5.2. Results
The PEST procedures are evaluated and weighed as stated in Section 4.1. Figure 12 shows the estimated IPD thresholds, as well as a representation of the change of the participant’s performance between the Org and Fov stimuli. The change rate is determined by dividing a participant’s threshold presented for an Org stimulus by the threshold of the respective Fov stimulus. This results in a positive change value in case the participant had a lower threshold in the Fov condition.
Neither the means nor the medians differ substantially for both the thresholds and change rates. The overlapping box plot notches yield a 95% certainty that the medians do not differ significantly, which is the case for all reported values. This suggests that participants exhibited similar stereoacuity for both rendering styles, and that the foveation did not compromise it. This finding is reflected by the constant change rate of ~0%, which shows that the participants’ performance did not change in either of the settings.
@@ 6. Future Work
It has been demonstrated that stereoacuity decreases in peripheral vision (Prince and Rogers, 1998; Mochizuki et al., 2012; Blakemore, 1970). Our results align with these findings and complement them by showing that this trend is not further enhanced by peripheral blur. This highlights the need for careful treatment of depth information in foveated content: Our findings indicate that reducing the spatial quality of an image does not entail that the quality of depth reproduction can be reduced as well. This observation is particularly important in applications that require depth processing. For example, stereoscopic and multiscopic content can be transmitted as a combination of rendered images and depth maps to recreate missing views during the decoding step (Merkle et al., 2009). When developing gaze-contingent compression methods, the compression of depth maps should be handled carefully and ideally separately from the compression of the image content. Similarly, any depth-related rendering process, such as displacement mapping or level-of-detail methods (Murphy and Duchowski, 2001; Surace et al., 2023), should be handled carefully and without implicitly assuming that foveation masks the imperfections in depth information. Exploring these issues in the future will broaden our understanding of the critical factors that must be replicated to achieve real-time realism in virtual and augmented realities.

### v2

@@ Abstract.
The true vision for real-time virtual and augmented reality is reproducing our visual reality in its entirety on immersive displays. To this end, foveated rendering leverages the limitations of spatial acuity in human peripheral vision to allocate computational resources to the fovea while reducing quality in the periphery. Such methods are often derived from studies on the spatial resolution of the human visual system and its ability to perceive blur in the periphery, enabling the potential for high spatial quality in real-time. However, the effects of blur on other visual cues that depend on luminance contrast, such as depth, remain largely unexplored. It is critical to understand this interplay, as accurate depth representation is a fundamental aspect of visual realism. In this paper, we present the first evaluation exploring the effects of foveated rendering on stereoscopic depth perception. We design a psychovisual experiment to quantitatively study the effects of peripheral blur on depth perception. Our analysis demonstrates that stereoscopic acuity remains unaffected (or even improves) by high levels of peripheral blur. Based on our studies, we derive a simple perceptual model that determines the amount of foveation that does not affect stereoacuity. Furthermore, we analyze the model in the context of common foveation practices reported in literature. The findings indicate that foveated rendering does not impact stereoscopic depth perception, and stereoacuity remains unaffected with up to 2×2$\times$2 × stronger foveation than commonly used. Finally, we conduct a validation experiment and show that our findings hold for complex natural stimuli.
@@
journalyear: 2025copyright: ccconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ; August 10–14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers (SIGGRAPH Conference Papers ’25), August 10–14, 2025, Vancouver, BC, Canadadoi: 10.1145/3721238.3730609isbn: 979-8-4007-1540-2/2025/08submissionid: 251ccs: Computing methodologies Perceptionccs: Computing methodologies Virtual reality
@@ 2.2. Depth Perception
To navigate 3D space, humans rely on a number of different depth cues that convey either relative or absolute distance information (Sweet and Kaiser, 2012; Cutting and Vishton, 1995). Cues such as accommodation, convergence, or familiar size convey information about the absolute distance between the observer and an object. Other cues provide solely relative information of ordinal or metric kind, allowing for the estimation of relative depth differences between objects. For example, occlusion conveys only ordinal information, whereas cues such as stereoscopic disparity or relative size provide metric information regarding the magnitude of depth differences.
With the advent of stereoscopic displays, the recreation of binocular vision became feasible. This enabled the display of stereo-
@@
scopic depth cues such as binocular disparity, resulting in stereopsis. Stereopsis is created by perceiving a difference in vergence angles between a focus point $p_{f}$ and a reference point $p_{r}$, which naturally occurs when those points are at different depths with respect to the eyes (Figure 2). The vergence angles can be described by $D(p)=\angle LpR$, denoting the angles between the two locations of the eyes in space ($L$ and $R$ ), and the respective points $p_{f}$ and $p_{r}$. Stereopsis is quantified by the disparity $d$ [], denoting the difference of these vergence angles:
(1) $d=D(p_{1})-D(p_{2}).$
The sign of $d$ denotes whether objects are relatively closer, or farther. In case both vergence angles are the same, $d=0$ and no disparity is present. This can happen when both points are at the same position in space, the image is rendered with an interpupillary distance (IPD) of 0mm, or both points lie on the horopter (Blakemore, 1970).
@@ 3. Methods
The stimulus design minimizes the influence of all depth cues except binocular disparity. While natural content presents a variety of depth cues such as motion parallax, shading, texture density, and deformation, the purpose of this study is to particularly measure disparity sensitivity in isolation. To this end, we ensure that monocular observation of the stimulus does not convey depth information, except for minor texture deformation. However, the visibility of the texture deformation is well below the resolvable disparity limit and, therefore, negligible (Bradshaw and Rogers, 1999).
@@ 3.1.1. Texture
50% Random Dot Patterns are widely used in vision science, particularly for researching depth perception (Julesz, 1960). They serve as a simple baseline pattern for our metric, exhibiting controllable features such as scale, frequency range, and dynamic range. As we aim to measure the lowest possible disparity threshold, we scale the random dot pattern to the optimally perceptible size. This scale equals the peak sensitivity value of the contrast sensitivity function (CSF) for the given eccentricity $\theta$. The stelaCSF [^1] https://github.com/gfxdisp/stelaCSF, retrieved 18. Oct 2024 (Mantiuk et al., 2022) yields a peak sensitivity $S_{F}(\theta)$ of $S_{F}(0^{\circ})=4.1$cpd, $S_{F}(10^{\circ})=1.8$cpd, and $S_{F}(20^{\circ})=1.3$cpd. We scale the dot pattern to match $S_{F}(\theta)$, by scaling the dots such that two dots equal one cycle.
@@
Additionally, the ring must be of a sufficient minimum width in the visual field to ensure good visibility. However, to reduce crosstalk between eccentricities and ensure accurate measurements, the width is minimized. To account for the varying sensitivity per eccentricity, we scale the ring wider with higher eccentricities. We set the ring’s width based on the CSF, choosing a width of 13 dots for both peripheral stimuli. This is approximately equal to $3.8^{\circ}$ in the visual field for $10^{\circ}$, and $5^{\circ}$ for $20^{\circ}$ eccentricity. For $0^{\circ}$ eccentricity we set a width of $6.7^{\circ}$, so that always 2 full cycles are visible. Additionally, the ring has a randomly dithered border (Figure 3). This attenuates the monoscopic depth cue of geometric deformation, which occurs due to the pixel shifting.
@@ 3.1.2. Depth
eye’s image, while leaving it untouched in the left eye’s image ($p_{L}$). This procedure follows implementations of previous work (Siderov and Harwerth, 1995; Bradshaw and Rogers, 1999; Didyk et al., 2011) and guarantees precise control of the presented disparity. The pixel shift introduces an intersection point $p_{I}$ which our brain perceives as being in front of the zero-plane (Figure 4). The disparity is measured between the point $p_{Z}$, which serves as a zero point reference (trough, $I_{D}(p_{Z})=0$), and $p_{I}$, the artificially created illusion point (peak, $I_{D}(p_{I})=1$). We generate the right eye’s image $I_{R}$ by resampling the left eye’s image $I_{L}$ at the respective image coordinates:
@@ 3.2.1. Procedure
Figure 7. Plot of estimated thresholds per eccentricity, as well as the polynomial fits. Each blue point denotes one measured threshold, the halo of each point denotes its weight $w$. The orange scatter plots denote the mean and standard error per $\sigma$. The line plots show the retrieved parabolas, fitted through the weighted threshold data.
[…]
@@ 3.2.3. Participants
The threshold estimation of $\theta=10^{\circ}$ and $\theta=20^{\circ}$ was conducted on 11 voluntary participants (1F, 10M), with an average age of 28 ± 4.7 years. All participants had normal or corrected-to-normal vision and were tested negative for stereo blindness. Nine participants reported prior experience in conducting perceptual experiments. The threshold estimation of $\theta=0^{\circ}$ was conducted on 10 of the 11 previous participants (1F, 9M; mean age 28.1 ± 4.9 years). Eight participants reported prior experience. With the exception of the two participating authors, all participants were naïve to the project. Additionally, all participants signed their previous consent and were compensated for their time. The study was approved by the institutional ethics committee.
@@ 4.1. Threshold Estimation
Since the PEST procedure can lead to sparse sampling of important regions and random answers can render data unreliable, the Weibull fit can get unstable. We obtain the final threshold $T$ as well as its standard deviation $T_{\sigma}$ by bootstrapping the data 100 times. This also yields a relative uncertainty of $u=\frac{T_{\sigma}}{T}$ per $T$. Measurements with $u>0.3$ are treated as outliers and excluded from further evaluation, as we consider the PEST procedure to be too unstable and therefore unreliable. This affects ~20% of our measurements. A plot of the excluded data points is provided in the Supplementary Material.
@@ 4.2. Perceptual Model
To estimate the parameters’ interpolation characteristics for varying $\theta$, we examine the characteristics of our fitted parabolas (Figure 7): The minimum threshold, associated with $p_{2}$, shifts to a higher $\sigma$ with increasing eccentricity. This indicates that the HVS tolerates higher blur levels in the periphery. Furthermore, the overall threshold $T$ shifts to higher values with increasing eccentricity, which resembles the y-offset encoded in $p_{3}$. The slope $p_{1}$ remains mostly unaltered. Figure 8 displays the parameter-values $p_{i}$ per eccentricity. As expected, they resemble the intuitive understanding discussed above. To model the parabolas for a continuous $\theta$, we fit an exponential curve of the form $p_{i}(\theta)=ae^{b\theta}+c$. An exponential curves increase strictly, which resembles our data. The fitting yields:
@@
This model holds for our measured scope of $\sigma\in[0,15]$, and $\theta\in[0,20]$. However, we note that $p_{1}$ exhibits a steep slope for higher eccentricities, resulting in extreme values for $\theta>20$. If such conditions are expected we suggest replacing $p_{1}$ with a constant fit of $p_{1}=0.0034$, which is the mean across $p_{1}(0^{\circ})$, $p_{1}(10^{\circ})$ and $p_{1}(20^{\circ})$. Replacing $p_{1}$ with this value only results in a minor difference to the current model, as the surface is still well-defined by $p_{2}$ and $p_{3}$, and $p_{1}$’s variability across the three measured eccentricities is not extreme.
@@ 5.1. Methods
We show participants half-split images of the natural scenes. The central part is blacked out up to $15^{\circ}$ eccentricity. We validate our model on high eccentricities, as the fovea remains almost unaffected by our model, yielding no measurable changes. As we allow the gaze to move within a margin of 3–5 from the fixation point, the effective blacking is ~10. Furthermore, an examination of $>15^{\circ}$ eccentricity retains enough binocular image space, as the headset only supports binocular vision up to $25^{\circ}$ eccentricity. The left and right halves of the image display different disparity levels: one half is rendered with 0mm IPD (no disparity), while the other half is rendered with a varying IPD (0–20mm) to control the overall magnitude of disparities in the scene (see Supplementary Material for vergence angle histograms). The IPD, and thus the overall scene depth, was adjusted in a staircase procedure. We simulate foveation by locally interpolating between pre-blurred images according to the per-pixel eccentricity value. Afterwards, we linearly map the values of the foveated image to match the value range of the input image.
Participants were shown both scenes in full-resolution (Org) and foveated (Fov) rendering styles. Each participant first viewed the Forest scene (F), and then the Kitchen scene (K). The order of presented rendering styles was randomized, and the style sequence was counterbalanced across scenes (e.g. F-Fov, F-Org, K-Org, K-Fov). Before threshold estimation began, participants were allowed to freely explore each scene using their central vision to familiarize themselves with prominent depth cues. This step was added to reduce measurement noise caused by initial variability in scene familiarity, as identified in pilot testing. Participants were asked to identify the side of the image which exhibited disparity. To estimate the IPD threshold, the displayed disparity was adjusted using the ”Best PEST” procedure, as described in Section 3.2.1. The experiment was conducted using the same apparatus as described in Section 3.2.2. We conducted the study on 15 participants (13M, 1F, 1X; mean age 26.1 ± 2.5 years). 11 participants reported prior experience with perceptual studies. All had normal or corrected-to-normal vision, signed their previous consent, and were compensated for their time. The study was approved by the institutional ethics committee.
@@ 5.2. Results
The PEST procedures are evaluated and weighed as stated in Section 4.1. Figure 12 shows the estimated IPD thresholds, as well as a representation of the change of the participant’s IPD threshold between the Org and Fov stimuli. The change rate is determined by dividing a participant’s threshold presented for an Org stimulus by the threshold of the respective Fov stimulus. This results in a positive change value in case the participant had a lower threshold in the Fov condition.
Neither the means nor the medians differ substantially for both the thresholds and change rates. The overlapping box plot notches yield a 95% certainty that the medians do not differ significantly, which is the case for all reported values. This suggests that participants exhibited similar stereoacuity for both rendering styles, and that the foveation did not compromise it. This finding is reflected by the constant change rate of ~0%, which shows that the participants’ performance did not change in either of the settings. Although the change rates vary up to ±$\pm 20\%$, we did not find a significant trend in personal preferences, meaning that participants did not have a similar change rate across conditions. In addition, the directions of change were not related to expert vs. inexperienced viewer.
@@ 6. Future Work
It has been demonstrated that stereoacuity decreases in peripheral vision (Prince and Rogers, 1998; Mochizuki et al., 2012; Blakemore, 1970). Our results align with these findings and complement them by showing that this trend is not further enhanced by peripheral blur. This highlights the need for careful treatment of depth information in foveated content: Our findings indicate that reducing the spatial quality of an image does not entail that the quality of depth reproduction can be reduced as well. This observation is particularly important in applications that require depth processing. For example, stereoscopic and multiscopic content can be transmitted as a combination of rendered images and depth maps to recreate missing views during the decoding step (Merkle et al., 2009). When developing gaze-contingent compression methods, the compression of depth maps should be handled carefully and ideally separately from the compression of the image content. Similarly, any depth-related rendering process, such as displacement mapping or level-of-detail methods (Murphy and Duchowski, 2001; Surace et al., 2023), should be handled carefully and without implicitly assuming that foveation masks naïve compression of depth information. Exploring these issues in the future will broaden our understanding of the critical factors that must be replicated to achieve real-time realism in virtual and augmented realities.
@@ 7. Conclusion
Acknowledgements.
This project has received funding from the European Research Council under the European Union’s Horizon 2020 research and innovation program (Grant 804226, PERDY).